# FIRE: A Failure-Adaptive Reinforcement Learning Framework for Edge Computing Migrations

### v3

@@ Abstract
Edge Computing, Service migration, Resilient Resource Allocation
@@ 1 Introduction
Mobile (or multi-access) edge computing (MEC) is a key technology in 5G and 6G telecommunication systems. It enables computationally intensive and latency sensitive mobile applications such as real time image processing, augmented reality, interactive gaming, etc. In MEC, computing resources such as server clusters are positioned close to end-users, typically at cellular base stations or WiFi access points at the edge of the radio access network [1, 2]. This proximity reduces response latency by avoiding the wide-area network delays encountered in cloud computing [2].
Researchers have been actively investigating how to jointly allocate computing and radio resources to computing tasks offloaded to edge servers [1, 3]. Nevertheless, user mobility across geographical areas presents a challenge towards task offloading [4, 5, 6]. Service migration, or moving service profiles to access points nearer the user as they move, has been proposed to maintain service continuity [5, 6]. Yet, frequent service migration causes additional expenditure, due to the required network usage. Therefore, balancing the delay-migration cost tradeoff has been widely studied [4, 7, 5, 8, 9]. Due to the large decision space when the numbers of access points and time-slots increase, and the lack of information on costs and mobility patterns, reinforcement learning (RL) is a common solution method [4, 7, 10]. Such methods often rely on digital twins of the edge computing network to train RL models, as digital twins allow the exploration of different resource allocation policies, without generating real-world consequences.
@@ 1.1 Challenges: Edge Computing and Resilience
Few migration works, however, account for another challenge in edge computing: the resilience of edge computing systems to rare but serious events like server failures. They occur due to reasons such as system overload, hardware failures, or malicious attacks, and they can be costly: the average cost of an outage at a cloud data center, for example, has increased to \$$\$740000$ in $2016$ [11]. Edge server failures may be even more probable than those in the cloud due to their distributed geographical locations, which complicates management and maintenance [12], though they will likely remain rare overall. Even worse, edge computing failures can have a sizeable impact as many edge computing applications are latency sensitive, such as the Internet of vehicles, augmented reality, and video processing applications [1]. Managing resources without considering rare but severe failure increases latency and reduces reliability. For instance, if the user’s service profile is migrated to a server that then fails, its job is not able to be completed, jeopardizing the smooth and safe functioning of applications, such as autonomous driving and real time obstacle detection.
We propose to use backups, which can take over when the primary service’s edge server fails, to handle failures. Nevertheless, managing the migration of both the primary and backup services is challenging, as the number of access points and timeslots increases. As mentioned above, due to the large decision space, and the network operator’s lack of information, RL has traditionally been proposed to solve the service migration problem [4, 7, 10, 13]. Nevertheless, rare events and failures occur at a low probability, making it difficult to jointly plan or learn an optimal resource allocation policy for both the usual and rare event scenarios. This is because reinforcement learning’s reliance on past reward data may overlook the significance of rare events, impacting the training process [14]. Therefore, we introduce FIRE: Failure-adaptive Importance sampling for Rare Events, a resilience framework for edge computing service migration.
@@ 1.2 FIRE: Failure-adaptive Importance sampling for Rare Events
The paper is structured as follows. Section 2 discusses the related work. Section 3 and 4 present the system models. Section 5 presents the problem formulation. Following which, we present our solution and its proofs in Section 6, and the deep Q-learning and Actor Critic algorithms in Section 7. Next, Section 8 addresses the heterogeneous risk tolerance scenario. Finally, we present simulation results in Section 9 and conclude in Section 10.
@@ 3 Migration of Individual Service Profiles
In this section, we introduce our failure-aware service migration model, with individual service profiles. In the next section, we will consider shared service profiles.
@@ 3.1 Service Placement Model
Users and edge servers. We consider an MEC system with $K$ users and $N$ access points (APs), such as a base station or wifi-access point. Each AP is equipped with a server to which users can offload latency-sensitive computing jobs, and we associate each AP with a disjoint physical region; users in this region then connect to this access point. Users are mobile and move from region to region at different times of the day. For example, office workers may move to the city for work during morning rush hours, changing their nearest AP. User $k$’s location at time $t$ is represented by $l_{u}^{k}(t)$. We consider that each user will be associated with its nearest AP, via local area network (LAN). The users’ mobility pattern follows a probability distribution $m(l_{u}^{\prime k}|l_{u}^{k})$, where $m(l_{u}^{\prime k}|l_{u}^{k})$ represents the transition probability the user travels to location $l_{u}^{\prime k}$, after being at location $l_{u}^{k}$.
Each user maintains a service profile at the edge servers, e.g., in virtual machines or containers [5], that consists of the state needed to run its edge computing application. Our goal is then to decide where the service profile and any backups should be located. To maintain low latencies for time-sensitive service applications, in light of user mobility, services are pre-migrated - likely to nearby APs. Service migration is stateful, involving the migration of the set of parameters and intermediate results associated with the user’s computing job. At each time-slot $t$, the user sends a service request to the network operator. The location of the user’s service at time $t$ is represented by $l_{s}^{k}(t)$, and due to pre-migrations the network operator makes this placement decision before the user location at time $t$ is known.
Failure model. In reality, rare events (system anomalies) such as server failures or shutdowns occur. Such events will impact the user’s quality of service, e.g., if the user is not able to have his or her job computed on time, impacting the safe and smooth functioning of edge applications and jeopardizing the low latencies of edge computing.
@@
Figure 1: Each user has its individual service profile, which has a backup, in case the server where there service profile is at fails.
@@
We formulate a Markov Decision Process (MDP) with system state $s(t)\in S$ is $[s_{1}(t),..s_{k}(t),...s_{K}(t)]$, where
@@
pre-emptively, before the users moves to their next locations $[l_{u}^{k}(t+1)]$. This involves determining the placement location of the service $l_{s}^{k}(t+1)$, and the position of the backup in the system $b_{u}^{k}(t+1)\in\{0,1,...N,\text{no backup}\}$. These decisions will directly impact the system state $s$ at the next timeslot.
@@
With a larger probability of $1-\epsilon(s)$, no server failure occurs, i.e. $f_{ind}=0$. The non rare event states “Normal States” are defined as the set $S\backslash T$. Hence we have
@@ 3.2.1 Operational Costs
The network operator migrates the users’ profiles across access points, to deliver a better QoS (lower job delay) for the user. We first model the costs incurred in normal (non-failure) states, which include communication and computing delays, both of which affect user QoS, as well as the cost of migrating service profiles. We let $d^{\text{comm},k}_{l_{u},j}$ denote the communication delay the user faces when it offloads its job to the edge server, given that its current location is $l_{u}^{k}$ and its service is placed at access point $j$. $d^{\text{comm},k}_{l_{u},j}$ consists of the access latency of uploading the job to the user’s associated access point, and the transfer latency of forwarding the job to the edge server’s location if the service is placed at another access point (i.e. $l_{u}^{k}\neq j$). This transferring latency via LAN depends on the hop count of the shortest communication path [30]. Therefore, the delay is a function of the distance between $l_{u}^{k}$ and $j$. Therefore, the total communication delay cost of $K$ users at time $t$ is
@@
where $\mathbbm{1}_{\{l_{s}^{k}=j\}}$ is the indicator function, which will take the value of $1$ if the user’s service profile is placed at access point $j$, and will take the value of 00 otherwise.
@@
where $\mathbbm{1}_{\{l_{s}^{k}(t-1)=i\}}$ is the indicator function that equals $1$ if the service profile was placed at location (AP) $i$ at time $t-1$.
@@ 3.2.2 Failure and Backup Costs
We next specify the costs incurred by provisioning backups in case of failures. There is a backup cost $B(t)$ incurred by storing the backup at a server: a sum of the backup migration cost and the backup storage cost, as it takes up space and prevents other content from being stored.
@@
where $\rho_{j}$ is the storage cost at access point $j$. There is also a failure cost $F(t)$ incurred
@@ 4 Migration of Shared Service Profiles
Figure 2: Multiple users share a service profile (SP) for their task, e.g. a common game environment or neural network. For instance, the users in red share the SP in red, which has multiple copies in the network.
In certain scenarios, multiple users share a common service profile (SP) for the same task. This could take the form of a common neural network for multi-modal learning, or a common game environment for their Augmented Reality (AR) or mobile gaming applications such as when multiple users are playing a game together [17]. Therefore, the network operator is concerned with the placement of shared SPs across the network, in light of the geographical demand distributions of the different SPs. In the event of a server failure, the impact will be multiplied when SPs are shared, as more users cannot have their job served.
As in the case of individual service profiles, we formulate a Markov Decision Process (MDP) in which the system state is $[s_{1,1}(t),..s_{i,j}(t),...s_{I,J}(t)]$, where for each location, SP pair ($i,j$), their component of the state space at time $t$ is
@@
where $d^{comm}_{i,\hat{i}}$ is the communication delay from location $i$ to $\hat{i}$, same as Section 3. The users at location $i$ access the nearest AP which contains the SP type they require ($\min_{\hat{i}\in\{\hat{i}|l^{SPj}_{\text{loc}\ \hat{i}}=1\}}d^{\text{comm}}_{i,\hat{i}}$). QoS is also affected by the total computing delay experienced across all APs at time $t$ [8]:
@@
The migration cost for SP $j$ kicks in if it is to be placed at location $i$ at time $t$ and was not placed there at $t-1$ (i.e. $d_{i,j}(t-1)=0\ \&\ d_{i,j}(t)=1)$. It consists of either downloading SP $j$ from the cloud server for a new deployment when $d_{\hat{i},j}(t-1)=0,\ \forall\hat{i}$, or migrating SP $j$ from the nearest location $\hat{i}$ when $\exists\hat{i}\ s.t.\ d_{\hat{i},j}(t)=1$. The total storage cost of storing all the deployed SPs is
@@
where $h_{j}$ is the size of service profile type $i$ and $\rho_{i}$ is the per unit storage cost at AP $i$. The failure cost is incurred for the users who need to use service profile type $j$ when either a) SP type $j$ is not placed anywhere ($l_{\text{loc}\ i}^{SPj}=0\ \forall i$), b) the delay exceeds $del^{j}$, the delay threshold for this application ($\min_{\hat{i}\in\{\hat{i}|l^{SPj}_{\text{loc}\ \hat{i}}=1\}}d^{\text{comm}}_{i,\hat{i}}>del^{j}\}$), or c) there is a failure at a location $\hat{i}$, and there is no available SP $j$ placed elsewhere, or the minimum delay elsewhere exceeds the delay threshold $del^{j}$ for this application.
@@ 5 Cost Minimization in Light of Rare Events
In this section, we introduce the service migration optimization problems, for both the individually used (Section 3) and shared (Section 4) service profile scenarios. Our focus is on long term service placements and backup storage decisions, taking into account potential rare events such as server shutdowns which have an impact on the user experience. For both problems, we aim to minimize the expected sum of the weighted sum of the communication and computing delay, migration, storage and failure costs. Actions are taken in a digital twin framework, as they allow for algorithm training while avoiding the cost of failures in real-world applications. Firstly, we introduce the optimization problem for the scenario where all users have their individual service profiles.
@@
where $w_{D},w_{C},$ etc. are the corresponding weights, which can be adjusted based on the priority of different costs. The decision variable $\pi(s,a)=\{P_{r}(a_{t}=a|s_{t}=s)\}$ is the network operator’s policy, the probability it will take each action $a(t)$ (service placement and backup storage decision for all users) given the state $s(t)$. The first constraint indicates that the state dynamics follows the transition probability of the system in Eq. (7), incorporating rare event transitions, while the other constraints indicate the delay, migration, backup and failure costs respectively. Next, we introduce the optimization problem for the shared service profiles scenario; note that we aim to minimize the same types of costs as in Eq. (18) with individual service profiles.
@@
Deciding a migration path for the user across $T$ timeslots is similar to the shortest path problem. However, the network operator may not have information on mobility distributions and the corresponding latency costs, complicating look-ahead solutions. Furthermore, the costs (e.g. computing delays, Eqs. (9) and (16)) of different users are coupled through service profiles’ sharing of AP resources. Thus, Markov Decision Processes and Reinforcement Learning have been used to derive service migration policies [7, 10, 13]. The difficulty in our problem further arises as we model occasional server failures. As these failures occur at a low probability, and because reinforcement learning relies on past reward data, the learner might overlook rare events and fail to learn an optimal failure-adaptive policy. We introduce our solution in the next section.
6 Reinforcement learning in the presence of Server Failures
In this section, we present FIRE: Failure-adaptive Importance sampling for Rare Events, a reinforcement learning based framework in an edge computing digital twin. While [16] and [14] also use importance sampling to sample rare events, these two works involve estimating the value function given a fixed policy. Their proof techniques also do not generalize to the $Q$-learning. In contrast, our algorithms involve learning the optimal policy $\pi^{*}$, which involves policy changes. Based on our algorithm, we propose novel boundedness and convergence proofs. We present our algorithm (FIRE-ImRE) and its proof in Subsections 6.2 and 6.3 respectively, and present an eligibility traces version (ETAA) for faster convergence in the Appendix. We propose a Deep Q learning and Actor Critic version of our algorithm in Section 7 to accommodate large state and action spaces. FIRE is compatible with any reinforcement learning formulation, and our algorithms are examples of this.

### v4

# FIRE: A Failure-Adaptive RL Framework for Edge Computing Migrations

@@ Abstract
Edge Computing, Service migration, Resilient Resource Allocation, Reinforcement Learning
@@ 1 Introduction
Mobile (or multi-access) edge computing (MEC) is a key technology in 5G and 6G telecommunication systems. It enables computationally intensive and latency sensitive mobile applications such as real time image processing, augmented reality (AR), interactive gaming, etc. In MEC, computing resources such as server clusters are positioned close to end-users, typically at cellular base stations or WiFi Access Points (AP) at the edge of the radio access network [1, 2]. This proximity reduces response latency by avoiding the wide-area network delays encountered in cloud computing [2].
Researchers have been actively investigating how to jointly allocate computing and radio resources to computing tasks offloaded to edge servers [1, 3]. Nevertheless, user mobility across geographical areas presents a challenge towards task offloading [4, 5, 6]. Service migration, or moving service profiles to APs nearer the user as they move, has been proposed to maintain service continuity [5, 6]. Yet, frequent service migration causes additional expenditure, due to the required network usage. Therefore, balancing the delay-migration cost tradeoff has been widely studied [4, 7, 5, 8, 9]. Due to the large decision space when the numbers of APs and time-slots increase, and the lack of information on costs and mobility patterns, Reinforcement Learning (RL) is a common solution method [4, 7, 10]. Such methods often rely on digital twins of the edge computing network to train RL models, as digital twins allow the exploration of different resource allocation policies, without generating real-world consequences.
@@ 1.1 Challenges: Edge Computing and Resilience
Few migration works, however, account for another challenge in edge computing: the resilience of edge computing systems to rare but serious events like server failures. They occur due to reasons such as system overload, hardware failures, or malicious attacks, and they can be costly: the average cost of an outage at a cloud data center, for example, has increased to $\mathdollar 740000$ in $2016$ [11]. Edge server failures may be even more probable than those in the cloud due to their distributed geographical locations, which complicates management and maintenance [12], though they will likely remain rare overall. Even worse, edge computing failures can have a sizeable impact as many edge computing applications are latency sensitive, such as the Internet of vehicles, AR, and video processing applications [1]. Managing resources without considering rare but severe failure increases latency and reduces reliability. For instance, if the user’s service profile is migrated to a server that then fails, its job is not able to be completed, jeopardizing the smooth and safe functioning of applications, such as autonomous driving and real time obstacle detection.
We propose to use backups, which can take over when the primary service’s edge server fails, to handle failures. Nevertheless, managing the migration of both the primary and backup services is challenging, as the number of APs and timeslots increases. As mentioned above, due to the large decision space, and the network operator’s lack of information, RL has traditionally been proposed to solve the service migration problem [4, 7, 10, 13]. Nevertheless, rare events and failures occur at a low probability, making it difficult to jointly plan or learn an optimal resource allocation policy for both the usual and rare event scenarios. This is because RL’s reliance on past reward data may overlook the significance of rare events, impacting the training process [14]. Therefore, we introduce FIRE: Failure-adaptive Importance sampling for Rare Events, a resilience framework for edge computing service migration.
@@ 1.2 FIRE: Failure-adaptive Importance sampling for Rare Events
The paper is structured as follows. Section 2 discusses the related work. Sections 3 and 4 present the system models, on the individual and shared service profile settings respectively. Section 5 presents the problem formulation. Following which, we present our solution and its proofs in Section 6, and the deep Q-learning and Actor Critic algorithms in Section 7. Next, Section 8 addresses the heterogeneous risk tolerance scenario. Finally, we present simulation results in Section 9 and conclude in Section 10.
@@ 2 Background and Related Work
We do not assume that failures can be predicted, unlike most other failure mitigation works. Instead, we aim to learn a policy that builds in resiliency to rare events that are not predicted in advance. Note that our algorithm can complement strategies like checkpointing [26] or dynamically adjusting the number of service backups. It addresses the unique challenge of resiliency in reinforcement learning-based policies, which fundamentally suffer from a lack of rare events in their training datasets.
@@ 3 Migration of Individual Service Profiles
In this section, we introduce our failure-aware service migration model, with individual service profiles. In the next section, we will consider shared service profiles. Our system models, which capture rare events, can be integrated into a “digital twin” for offline algorithm learning in light of rare events such as server failures. The purpose of digital twins is to allow the system to learn a policy in a simulated and virtual environment, without actually experiencing the cost of rare events. For example, a simulated edge computing environment may serve as a digital twin [31, 32].
@@ 3.1 Service Placement Model
Users and edge servers. We consider an MEC system with $K$ users and $N$ APs (APs), such as a base station or wifi-access point. Each AP is equipped with a server to which users can offload latency-sensitive computing jobs, and we associate each AP with a disjoint physical region; users in this region then connect to this AP. Users are mobile and move from region to region at different times of the day. For example, office workers may move to the city for work during morning rush hours, changing their nearest AP. User $k$’s location at time $t$ is represented by $l_{u}^{k}(t)$. We consider that each user will be associated with its nearest AP, via local area network (LAN). The users’ mobility pattern follows a probability distribution $m(l_{u}^{\prime k}|l_{u}^{k})$, where $m(l_{u}^{\prime k}|l_{u}^{k})$ represents the transition probability the user travels to location $l_{u}^{\prime k}$, after being at location $l_{u}^{k}$.
Each user maintains a service profile at the edge servers, e.g., in VMs or containers [5], that consists of the state needed to run its edge computing application. Our goal is then to decide where the service profile and any backups should be located. To maintain low latencies for time-sensitive service applications, in light of user mobility, services are pre-migrated - likely to nearby APs. Service migration is stateful, involving the migration of the set of parameters and intermediate results associated with the user’s computing job. At each time-slot $t$, the user sends a service request to the network operator. The location of the user’s service at time $t$ is represented by $l_{s}^{k}(t)$, and due to pre-migrations the network operator makes this placement decision before the user location at time $t$ is known.
Failure model. In reality, rare events (system anomalies) such as server failures or shutdowns occur. Such events will impact the user’s quality of service, e.g., if the user is not able to have his or her job computed on time, impacting the safe and smooth functioning of edge applications and jeopardizing the low latencies of edge computing. We define a failure as any event that prevents a user’s service job from being completed, which may stem from either a hardware or a software failure.
@@
Figure 1: Users move across the network. Each user has its individual service profile, which has a backup, in case the server where their service profile is at fails. To evaluate system resilience and optimize failure-handling algorithms and strategies without incurring real-world costs, the setup is modeled within a digital twin.
@@
We formulate a Markov Decision Process (MDP) with system state $s(t)\in S$ is $s_{1}(t),..s_{k}(t),...s_{K}(t)]$, where
@@
pre-emptively, before the users moves to their next locations $l_{u}^{k}(t+1)]$. This involves determining the placement location of the service $l_{s}^{k}(t+1)$, and the position of the backup in the system $b_{u}^{k}(t+1)\in\{0,1,...N,\text{no backup}\}$. These decisions will directly impact the system state $s$ at the next timeslot.
@@
[…]
@@ 3.2.1 Operational Costs
The network operator migrates the users’ profiles across APs, to deliver a better QoS (lower job delay) for the user. We first model the costs incurred in normal (non-failure) states, which include communication and computing delays, both of which affect user QoS, as well as the cost of migrating service profiles. We let $d^{\text{comm},k}_{l_{u},j}$ denote the communication delay the user faces when it offloads its job to the edge server, given that its current location is $l_{u}^{k}$ and its service is placed at AP $j$. $d^{\text{comm},k}_{l_{u},j}$ consists of the access latency of uploading the job to the user’s associated AP, and the transfer latency of forwarding the job to the edge server’s location if the service is placed at another AP (i.e. $l_{u}^{k}\neq j$). This transferring latency via LAN depends on the hop count of the shortest communication path [33]. Therefore, the delay is a function of the distance between $l_{u}^{k}$ and $j$. Therefore, the total communication delay cost of $K$ users at time $t$ is
@@
where $\mathbbm{1}_{\{l_{s}^{k}=j\}}$ is the indicator function, which will take the value of $1$ if the user’s service profile is placed at AP $j$, and will take the value of 0 otherwise.
@@
where $\mathbbm{1}_{\{l_{s}^{k}(t-1)=i\}}$ is the indicator function that equals $1$ if the service profile was placed at location (AP) $i$ at time t1t-1.
@@ 3.2.2 Failure and Backup Costs
We next specify the failure costs and the costs incurred by provisioning backups in case of failures. As mentioned earlier, the system can be modeled within a digital twin setup, allowing algorithm testing and policy learning in a virtual and simulated setting - to avoid experiencing the real-world cost. There is a backup cost $B(t)$ incurred by storing the backup at a server: a sum of the backup migration cost and the backup storage cost, as it takes up space and prevents other content from being stored.
@@
where $\rho_{j}$ is the storage cost at AP $j$. There is also a failure cost $F(t)$ incurred
@@ 4 Migration of Shared Service Profiles
Figure 2: Multiple users (who are mobile across the network) share a service profile (SP) for their task, e.g. a common game environment or neural network. For instance, the users in red share the SP in red, which has multiple copies in the network.
In this section, we introduce the edge computing service migration scenario where, differing from Section 3, multiple users share a common service profile (SP) for the same task. This could be a common neural network for multi-modal learning, or a common game environment for their AR or mobile gaming applications such as when multiple users are playing a game together [17]. Therefore, the network operator is concerned with the placement of shared SPs across the network, in light of the geographical demand distributions of the different SPs. In the event of a server failure, the impact will be multiplied when SPs are shared, as more users cannot have their job served.
As in the case of individual service profiles, we formulate a Markov Decision Process (MDP) in which the system state is $s_{1,1}(t),..s_{i,j}(t),...s_{I,J}(t)]$, where for each location, SP pair ($i,j$), their component of the state space at time $t$ is
@@
[…]
@@
The migration cost for SP $j$ kicks in if it is to be placed at location $i$ at time $t$ and was not placed there at t1t-1 (i.e. $d_{i,j}(t-1)=0\ \&\ d_{i,j}(t)=1)$. It consists of either downloading SP $j$ from the cloud server for a new deployment when $d_{\hat{i},j}(t-1)=0,\ \forall\hat{i}$, or migrating SP $j$ from the nearest location $\hat{i}$ when $\exists\hat{i}\ s.t.\ d_{\hat{i},j}(t)=1$. The total storage cost of storing all the deployed SPs is
@@
[…]
@@ 5 Cost Minimization in Light of Rare Events
In this section, we introduce the service migration optimization problems, for both the individually used (Section 3) and shared (Section 4) service profile scenarios. Our focus is on long term service placements and backup storage decisions, taking into account potential rare events such as server shutdowns ($f^{k}_{ind}(t)=1$) which have an impact on the user experience. For both problems, we aim to minimize the expected sum of the weighted sum of the communication and computing delay, migration, storage and failure costs. Actions are taken in a digital twin framework, as they allow for algorithm training while avoiding the cost of failures in real-world applications. Firstly, we introduce the optimization problem for the scenario in Section 3 where all users have their individual service profiles.
@@
where $w_{D},w_{C},$ etc. are the corresponding weights, which can be adjusted based on the priority of different costs. The decision variable $\pi(s,a)=\{P_{r}(a_{t}=a|s_{t}=s)\}$ is the network operator’s policy, the probability it will take each action $a(t)$ (service placement and backup storage decision for all users) given the state $s(t)$. The first constraint indicates that the state dynamics follows the transition probability of the system in Eq. (7), incorporating rare event transitions, while the other constraints indicate the delay, migration, backup and failure costs respectively.
Next, we introduce the optimization problem for the scenario in Section 4 where users share service profiles; note that we aim to minimize the same types of costs as in Eq. (18) with individual service profiles.
@@
Deciding a migration path for the user across $T$ timeslots is similar to the shortest path problem. However, the network operator may not have information on mobility distributions and the corresponding latency costs, complicating look-ahead solutions. Furthermore, the costs (e.g. computing delays, Eqs. (9) and (16)) of different users are coupled through service profiles’ sharing of AP resources. Thus, Markov Decision Processes and RL have been used to derive service migration policies [7, 10, 13]. The difficulty in our problem further arises as we model occasional server failures. As these failures occur at a low probability, and because RL relies on past reward data, the learner might overlook rare events and fail to learn an optimal failure-adaptive policy. We introduce our solution in the next section.
6 RL in the presence of Server Failures
In this section, we present FIRE: Failure-adaptive Importance sampling for Rare Events, a RL based framework in an edge computing digital twin. Our algorithm is applied to solve both optimization problems introduced in Section 5. While [16] and [14] also use importance sampling to sample rare events, these two works involve estimating the value function given a fixed policy. Their proof techniques also do not generalize to the $Q$-learning. In contrast, our algorithms involve learning the optimal policy $\pi^{*}$, which involves policy changes. Based on our algorithm, we propose novel boundedness and convergence proofs. We present our algorithm (FIRE-ImRE) and its proof in Subsections 6.2 and 6.3 respectively, and present an eligibility traces version (ETAA) for faster convergence in the Appendix. We propose a Deep Q learning and Actor Critic version of our algorithm in Section 7 to accommodate large state and action spaces. FIRE is compatible with any RL formulation, and our algorithms are examples of this.